# From Concept to Measurement: A Survey of How the Blockchain Trilemma Can Be Analyzed

### v2

@@ II-A Blockchain Technology
Blockchain technology is a type of distributed ledger technology that enables the operation of blockchain systems [21]. Blockchain systems store replications of a ledger (i.e., a record of transactions) on distributed computing devices called nodes. The ledger is structured into blocks, each containing a batch of transactions. Except for the genesis block, each block references its preceding block with that predecessor’s hash value [22]. A sequence of blocks, each referencing its predecessor, forms a blockchain.
@@
Most blockchain systems (e.g., the Bitcoin and Ethereum systems) operate as replicated state machines, where validating nodes maintain local states by redundantly executing the same protocol (see Figure 1), such as for validating and verifying transactions and blocks. The local state of a validating node is defined by its stored blockchain.
@@
To publish a new transaction, a validating node broadcasts the transaction to its adjacent validating nodes. Upon receipt, validating node processes incoming transactions redundantly by executing the same protocol. Upon successfully validating a transaction, a validating node’s state transitions to a subsequent state. In this work, we focus on validating nodes, excluding clients, as highlighted by the dashed border in Figure 1.
@@
Validating nodes use consensus mechanisms to reach consistency across all replicas. Consensus mechanisms define key properties of blockchain systems, including their permission models, finality models, and fault tolerance. In terms of permission models, a blockchain system can be permissionless, where all validating nodes have equal rights to participate (e.g., in the Bitcoin system), or permissioned, where participation is restricted based on predefined rules [24, 21], such as in systems using the Hyperledger Fabric protocol. Such permissions determine who can join the network and participate in consensus finding.
Finality models define when blocks appended to the blockchain are considered finalized, meaning they cannot be altered or reverted. In immediate (or deterministic) finality, a block is considered finalized as soon as it is added to the blockchain. In contrast, probabilistic finality assumes that a block becomes increasingly final as more successor blocks are appended to it, decreasing the likelihood of reorganization over time.
In terms of fault tolerance, blockchain systems can be omission-tolerant, crash-fault–tolerant, and Byzantine fault–tolerant [21, 1, 25, 26]. Omission tolerance refers to blockchain systems that can compensate for network messages that are lost in transit. Crash-fault tolerance refers to the ability of a blockchain system to compensate for validating nodes that are (temporarily) unavailable, for example, because the validating nodes crashed or the network connection is unreliable. Byzantine fault tolerance [25, 26] extends crash-fault tolerance by the ability to compensate for accidental faults and deliberate attacks. Accidental faults include software bugs and misconfiguration, while adversarial attacks involve strategies, such as in selfish mining [27, 28, 29].
Consensus mechanisms strongly influence performance of blockchain systems [21, 1], such as in terms of transaction processing rates (i.e., throughput). Voting-based consensus mechanisms with immediate finality, such as Practical Byzantine fault tolerance (PBFT) [30, 31], experience performance degradation as network size increases. This is primarily due to the higher communication complexity required for consensus among an increasing number of validating nodes. For consensus mechanisms with probabilistic finality, such as Nakamoto consensus in the Bitcoin system [22], throughput is often less affected by changes in network size, but consistency assumptions are strongly relaxed. Instead of ensuring that all validating nodes store the same version of the ledger at all times, they provide eventual consistency, where most validating nodes converge to a consistent state over time [21].
@@ IV-A1 Degree of Decentralization
$Gini=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}\left|t_{i}-t_{j}\right|}{2\times n}$ (12)
@@
Equation 12 can be used as follows. Suppose five validating nodes ${n_{0},n_{1},n_{2},n_{3},n_{4},}$ own the following tokens: $n_{0}=2$, $n_{1}=3$, $n_{2}=5$, $n_{3}=6$, and $n_{4}=2$. The mean differences of tokens owned by validating nodes can be calculated in pairs as follows:
@@
$Gini=\frac{8+7+9+12+8}{2\times 5}=4.40$
If a single validating node owned 18 tokens while other validating nodes owned none, the $Gini$ coefficient equals 14.40, indicating a lower DoD compared to 4.40 shown in the previous example. If all validating nodes owned equal amounts of tokens $Gini=0$. The value of the $Gini$ coefficient obtained in the earlier example (i.e., 4.4) indicates that validating nodes have fairly equitable wealth, indicating a higher DoD.
The $Gini$ coefficient offers a simple estimate of the DoD based on wealth distribution to validating nodes. The metric, however, underrepresents other considerations (e.g., use of reputation in weighted PoS) [72] sometimes used in consensus finding in addition to token ownership. Amounts of tokens owned by validating, therefore, do not guarantee their actual participation in consensus finding. Moreover, the metric neglects other factors that can influence consensus in addition to wealth distribution (e.g., bandwidth). Validating nodes with high bandwidth often have more influence on consensus finding than those with low bandwidth [1, 61].
@@ IV-A2 Scalability
Equations 21 and 22 apply only to throughput, neglecting scalability in terms of the number of validating nodes in blockchain systems. The number of validating nodes, however, influences communication complexity in consensus finding, thus the time until transactions are confirmed [85]. Moreover, the metrics only focus on confirmed transactions, neglecting overhead to process transactions that are not included in the mainchain. Neglecting such transactions may cause equations 21 to underestimate maximum throughput.
@@ IV-A3 Security
$ConfLat_{F}$ denotes the confirmation latency when a blockchain system experiences faults. Input variables in equations 32 and equations 33 can be obtained using equations 21, 22, 15, and 16.
@@ Appendix A Overview of Input Variables
Block Confirmation Time The timestamp in milliseconds when a new block is (assumed to be) confirmed in a blockchain system. 15, 29
Block Creation Interval The average time in milliseconds between the proposal of consecutive blocks that are included in a blockchain. 16, 22
Block Gas Cost The computation cost required to execute a transaction in a blockchain system. 22
Block Gas Limit The maximum amount of gas available in a block to process a transaction in a blockchain system. 22
Block Proposal Time The timestamp in milliseconds when a new block is proposed to a blockchain system. 15, 29
Hashing Power The number of hashes per second used to produce a new block. 9, 30
MemPool Size The maximum possible number of pending transactions buffered in a mempool. 22
Number of Confirmed Blocks The number of (probabilistically) finalized blocks stored in a blockchain system. 29, 35
@@
Transactions The total number of transactions processed and included in a block that has been (probabilistically) finalized into the mainchain of a blockchain system. 14, 21
Number of Failures The number of failures in a blockchain system within a given observation timespan. 26
@@
Number of Repairs The number of repairs, including recovery, a blockchain system experiences after failures within a defined timespan. 26
Number of Stale Blocks The number of valid blocks that are proposed but eventually not included in the mainchain. 35
Number of Tokens The number of tokens owned by an individual validating node. 10, 12, 30
Number of Transactions The total number of transactions issued to a blockchain system. 14
Number of Validating Nodes The number of validating nodes in a blockchain system. 1, 3, 10, 9, 30, 12
Resource Cost The value of tokens or hashing power in fiat currency. 30
Security Confirmation The required minimum number of blocks that must be appended to a specific block to achieve sufficiently high probabilistic finality for that block. 16
@@
Total Operational Time The timespan a blockchain system operates correctly within a define observation time. 26, 34
Total Repair Time The timespan a blockchain system takes to recover from failure. 26

### v3

@@ II-A Blockchain Technology
Blockchain technology is a special form of distributed ledger technology that enables the operation of blockchain systems—distributed databases designed to record transactions securely and consistently [21]. Blockchain systems store replicas of a record of transactions on distributed computing devices called nodes. The record is structured into blocks, each containing a batch of transactions. Except for the genesis block, each block references its preceding block with that predecessor’s hash value [22]. A sequence of blocks, each referencing its predecessor, forms a blockchain.
@@
Most blockchain systems (e.g., the Bitcoin and Ethereum systems) operate as replicated state machines, where validating nodes maintain consistent local replicas of the record by redundantly executing the same protocol (see Figure 1), such as for validating and verifying transactions and blocks. The local state of a validating node is defined by its stored blockchain.
@@
When a new transaction is created, a node broadcasts it to its adjacent validating nodes. Each validating node processes incoming transactions by executing the same protocol, ensuring redundant validation. Upon successful validation, the node’s state transitions accordingly.
To maintain consistency among replicas, validating nodes use consensus mechanisms. Consensus mechanisms shape key properties of blockchain systems, including their permission models, finality models, and fault tolerance. In terms of permission models, a blockchain system can be permissionless, where all validating nodes have equal rights to participate (e.g., in the Bitcoin system), or permissioned, where participation is restricted based on predefined rules [23, 21], such as in systems using the Hyperledger Fabric protocol. Such permissions determine who can join the network and participate in consensus finding.
Finality models define when blocks appended to the blockchain are considered finalized, meaning they cannot be altered or reverted. In immediate (or deterministic) finality, a block is considered finalized as soon as it is added to the blockchain. In probabilistic finality, a block’s finality increases as more blocks are added on top, reducing the likelihood of reorganization over time.
In terms of fault tolerance, blockchain systems can be omission-tolerant, crash fault-tolerant, and Byzantine fault-tolerant [21, 1, 24, 25]. Omission tolerance refers to blockchain systems that can compensate for network messages that are lost in transit. Crash-fault tolerance refers to the ability of a blockchain system to compensate for validating nodes that are (temporarily) unavailable, for example, because the validating nodes crashed or the network connection is unreliable. Byzantine fault tolerance [24, 25] extends crash-fault tolerance by the ability to compensate for accidental faults and deliberate attacks. Accidental faults include software bugs and misconfiguration, while adversarial attacks involve strategies, such as in selfish mining [26, 27, 28].
Consensus mechanisms strongly influence the performance of blockchain systems [21, 1], especially in terms of transaction processing rates (i.e., throughput). Voting-based consensus mechanisms with immediate finality, such as Practical Byzantine fault tolerance (PBFT) [29, 30], experience performance degradation as network size increases. This is mainly due to the higher communication complexity required for consensus among an increasing number of validating nodes. For consensus mechanisms with probabilistic finality, such as Nakamoto consensus in the Bitcoin system [22], throughput is often less affected by changes in network size, but consistency assumptions are strongly relaxed. Instead of ensuring that all validating nodes store the same version of the replica at all times, they provide eventual consistency, where most validating nodes converge to a consistent state over time [21].
While blockchain technology [22] is a prominent example of distributed ledger technology, the underlying principles, such as consensus and replicated state machines, are not unique to blockchains. Many distributed systems achieve consistency and fault tolerance through similar mechanisms, even if they do not employ a strict blockchain data structure. For instance, Hyperledger Fabric utilizes traditional databases to store data, while IOTA organizes transactions in a directed acyclic graph (DAG) rather than a linear sequence of blocks.
Although the ‘blockchain trilemma’ was originally formulated to describe the trade-offs faced by blockchain systems, similar tensions arise in a broader class of distributed databases that rely on replication and consensus. Thus, the blockchain trilemma encapsulates a fundamental design challenge: achieving an optimal balance among DoD, scalability, and security is inherently difficult in various distributed database systems that seek consensus across nodes. This observation extends the relevance of the trilemma beyond blockchain systems to distributed databases that use replicated state machines and consensus. This work considers such distributed databases with a focus on blockchain systems. It focuses on the role of validating nodes or peers in general distributed databases, excluding client nodes, as indicated by the dashed border in Figure 1.
@@ IV-A1 Degree of Decentralization
$Gini=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}\left|t_{i}-t_{j}\right|}{2n\sum_{i=1}^{n}t_{i}}$ (12)
@@
Equation 12 can be used as follows. Suppose five validating nodes ${n_{1},n_{2},n_{3},n_{4},n_{5},}$ own the following tokens: $n_{1}=2$, $n_{2}=3$, $n_{3}=5$, $n_{4}=6$, and $n_{5}=2$. The mean differences of tokens owned by validating nodes can be calculated in pairs as follows:
@@
$Gini=\frac{8+7+9+12+8}{2\times 5\times 18}\approx 0.244$
If a single validating node owned 18 tokens while other validating nodes owned none, the $Gini$ coefficient equals 0.8, indicating a lower DoD compared to 0.244 shown in the previous example. If all validating nodes owned equal amounts of tokens $Gini=0$. The value of the $Gini$ coefficient obtained in the earlier example (i.e., 0.244) indicates that validating nodes have fairly equitable wealth, indicating a higher DoD.
The $Gini$ coefficient is an economic metric that offers a simple estimate of the DoD based on wealth distribution. The metric, however, underrepresents other considerations (e.g., use of reputation in weighted PoS) [72] sometimes used in consensus finding in addition to token ownership. Amounts of tokens owned by validating, therefore, do not guarantee their actual participation in consensus finding. Moreover, the metric neglects other factors that can influence consensus in addition to wealth distribution (e.g., bandwidth). Validating nodes with high bandwidth often have more influence on consensus finding than those with low bandwidth [1, 61]. Additionally, for the metric, it is assumed that each validating node is independently controlled. In practice, however, a single entity may operate multiple nodes and hold their tokens. To correctly use $Gini$ coefficient, it is essential to account for such social factors.
@@ IV-A2 Scalability
Equations 22 and 23 apply only to throughput, neglecting scalability in terms of the number of validating nodes in blockchain systems. The number of validating nodes, however, influences communication complexity in consensus finding, thus the time until transactions are confirmed [85]. Moreover, the metrics only focus on confirmed transactions, neglecting overhead to process transactions that are not included in the mainchain. Neglecting such transactions may cause equations 22 to underestimate maximum throughput.
@@ IV-A3 Security
$ConfLat_{F}$ denotes the confirmation latency when a blockchain system experiences faults. Input variables in equations 33 and equations 34 can be obtained using equations 22, 23, 16, and 17.
@@ Appendix A Overview of Input Variables
Block Confirmation Time The timestamp in milliseconds when a new block is (assumed to be) confirmed in a blockchain system. 16, 30
Block Creation Interval The average time in milliseconds between the proposal of consecutive blocks that are included in a blockchain. 17, 23
Block Gas Cost The computation cost required to execute a transaction in a blockchain system. 23
Block Gas Limit The maximum amount of gas available in a block to process a transaction in a blockchain system. 23
Block Proposal Time The timestamp in milliseconds when a new block is proposed to a blockchain system. 16, 30
Hashing Power The number of hashes per second used to produce a new block. 9, 31
MemPool Size The maximum possible number of pending transactions buffered in a mempool. 23
Number of Confirmed Blocks The number of (probabilistically) finalized blocks stored in a blockchain system. 30, 36
@@
Transactions The total number of transactions processed and included in a block that has been (probabilistically) finalized into the mainchain of a blockchain system. 15, 22
Number of Failures The number of failures in a blockchain system within a given observation timespan. 27
@@
Number of Repairs The number of repairs, including recovery, a blockchain system experiences after failures within a defined timespan. 27
Number of Stale Blocks The number of valid blocks that are proposed but eventually not included in the mainchain. 36
Number of Tokens The number of tokens owned by an individual validating node. 10, 12, 31
Number of Transactions The total number of transactions issued to a blockchain system. 15
Number of Validating Nodes The number of validating nodes in a blockchain system. 1, 3, 10, 9, 31, 12
Resource Cost The value of tokens or hashing power in fiat currency. 31
Security Confirmation The required minimum number of blocks that must be appended to a specific block to achieve sufficiently high probabilistic finality for that block. 17
@@
Total Operational Time The timespan a blockchain system operates correctly within a define observation time. 27, 35
Total Repair Time The timespan a blockchain system takes to recover from failure. 27